OpenStudy (anonymous):

I have math problems pictures, is anyone willing to swap emails to help me, im on mobile device which wont let me draw. I really do need help badly. Im desperate.

OpenStudy (anonymous):

Pleaseeee,

OpenStudy (omarbirjas):

Hello @hav5v15estjohn , you must be new here, unfortunately we cannot help you by sending you emails for security reasons take the time to read the Code of Conduct, the Terms and Conditions, and the Privacy Policy bellow: http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions http://openstudy.com/privacy-policy However, I'll gladly try to help you on Openstudy.

OpenStudy (anonymous):

Well most of my problems are triangles. Unfortunately icannot draw on this because of my phone

OpenStudy (anonymous):

Is there a way youcan add an image using a cell phone?f

OpenStudy (omarbirjas):

You can screen shot it, and attach the file, or you can go to http://prnt.sc/ and try uploading your image there, then send me the link that it gives you.

OpenStudy (anonymous):

http://prntscr.com/79epal

OpenStudy (anonymous):

http://prntscr.com/79eqvt Im not good at math at all.

OpenStudy (omarbirjas):

The formula to find the area of any parallelogram is a*h=Area a being the "unslanted" edge, and h being the height. As you can see here we have ben given an acute angle, a and a "slanted" side.

OpenStudy (omarbirjas):

To find the length of the height we are going to use a little bit of trigonometry.

OpenStudy (omarbirjas):

\[\sin \theta \frac{ opposite }{ hypotenuse }\]

OpenStudy (omarbirjas):

Plug in the information we have and you get \[\sin 60 \frac{ d }{ 20 }\]

OpenStudy (anonymous):

I wrote all of that down. ForThe firsT problem

OpenStudy (omarbirjas):

Sorry, put an equal sign between all the sin 60s and d/20 etc.

OpenStudy (omarbirjas):

What problem do you need help with?

OpenStudy (anonymous):

I need help with every one ofthem. t

OpenStudy (omarbirjas):

Ok then, back to where I was, plugging that in will give you that sin 60 =0.8660

OpenStudy (omarbirjas):

0.8660=x/20

OpenStudy (omarbirjas):

x=17.32

OpenStudy (omarbirjas):

Now we have the height, it's all smooth sailing from here. What do you do to find the area of a parallelogram?

OpenStudy (anonymous):

The permiter and area

OpenStudy (omarbirjas):

Ok, we need to find those? But do you know how? Lets start with the perimeter. How do you find the perimeter.

OpenStudy (omarbirjas):

Sorry for my punctuation.

OpenStudy (anonymous):

I have no idea . How to find them.

OpenStudy (omarbirjas):

Really? You're doing trigonometry and you can't find a simple perimeter of a parallelogram? It's ok, no biggie, I'll show you.

OpenStudy (anonymous):

I was never good at math. Lol

OpenStudy (omarbirjas):

Perimeter is the the length of all the sides combined, when know that the "slanted" sides have a length of 20 each and the "straight" sides have a length of 24 each. Finding the perimeter, all we have to do is add 24+24+20+20= what do you get?

OpenStudy (omarbirjas):

88

OpenStudy (omarbirjas):

Now for area, all you have to do is multiply height by the perpendicular side, or the "straight" side.

OpenStudy (omarbirjas):

17.32*24= 415.68

OpenStudy (omarbirjas):

round 415.68 and you get 415.7

OpenStudy (omarbirjas):

Your answer is D.

OpenStudy (omarbirjas):

Now for the second one, 6*3.4= area 6+6+3.4+3.4= perimeter

OpenStudy (anonymous):

Thank you. Ijust need help with the .

OpenStudy (anonymous):

Rest

OpenStudy (omarbirjas):

No problem, if you ever need any help with navigating openstudy, fan me and give me a testimonial telling me to fan you, I'll PM you anything you need.

OpenStudy (omarbirjas):

As for the rest, I can't help you, try tagging users with the @ symbol and putting their user name like this @ha5v15estjohn

OpenStudy (omarbirjas):

I'm very busy, my apologies.

OpenStudy (anonymous):

Thank You